OpenStudy (anonymous):

Choose the correct simplification of (a3b2)(ab). a3b2 a5b5 a6b6 a4b3

OpenStudy (cwrw238):

a^3 * a = a^4 b^2 * b = b^3

OpenStudy (anonymous):

(a3b2)(ab) a3+1 b2+1 a4 b3 So it's the last option :)

OpenStudy (cwrw238):

yea